Circular No. 6677

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 L1 On June 5, J. Zhu, Beijing Astronomical Observatory, reported ob-

servations of five new objects detected on June 3 and 4 in the course of the BAO Schmidt CCD Asteroid Program at Xinglong. While observing some of these objects on June 8, D. D. Balam, Victoria, noted that one was clearly cometary, something that was confirmed by observers in Europe. CCD precise positions:

1997 UT α_{2000} δ_{2000} m_{1} Observer
June 3.67351 15^{h}31^{m}55.^{s}45 −6^{o}30′17′′.3  Zhu
3.68338 15 31 55.21 −6 30 10.0  _{”}
3.69281 15 31 54.85 −6 30 04.9 16.7  _{”}
4.68358 15 31 21.55 −6 20 40.0 17.1  _{”}
4.69367 15 31 21.18 −6 20 35.2  _{”}
4.70343 15 31 20.81 −6 20 29.3  _{”}
8.36119 15 29 21.84 −5 46 40.9 16.5 Balam
8.37000 15 29 21.56 −5 46 35.8 16.5  _{”}
8.85532 15 29 06.37 −5 42 14.4  Vagnozzi
8.88558 15 29 05.39 −5 41 58.1 17.8 Tichá
8.89414 15 29 05.13 −5 41 52.1  _{”}
8.89819 15 29 05.02 −5 41 52.1  Vagnozzi

J. Zhu, Y. J. Chen and X. Y. Li (Xinglong). 0.6-m Schmidt. Measurers

C. M. Ma, J. Zhu and X. Y. Li. Nothing reported about appearance. D. D. Balam (Dominion Astrophysical Observatory). 1.8-m reflector. Ob-

ject diffuse, tail 1′ long in p.a. 220^{o}. A. Vagnozzi (Stroncone). 0.5-m f/2.8 Ritchey-Chrétien. A longer exposure

shows a tail 1′.43 long in p.a. 205^{o} deg. J. Tichá and M. Tichỳ (Kleť). 0.57-m f/5.2 reflector. Slightly diffuse.

Parabolic orbital elements by B. G. Marsden from the above observations:

T = 1998 Jan. 24.330 TT  ω = 83.^{o}677
Ω = 224.946 }2000.0
q = 2.39292 AU  i = 33.906
1997TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 6 15^{h}30.^{m}63 − 6^{o}08.^{′}3 2.514 3.452 153.^{o}5 7.^{o}5 16.9
11 15 28.01 − 5 23.1 2.510 3.416 148.3  9.0 16.8
16 15 25.67 − 4 40.9 2.513 3.380 143.1 10.4 16.8
21 15 23.66 − 4 02.0 2.521 3.345 137.9 11.8 16.8
1997 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden